OpenStudy (anonymous):

a 20 foot ladder rests against a building 15 feet from the floor. how far does the ladder extend from the base of the wall? what angle does the ladder make with the ground?

OpenStudy (anonymous):

I believe this is the scenario , yes?

OpenStudy (anonymous):

yes that's what i drew . i'm not sure how to get the angle from there

OpenStudy (anonymous):

you could take the arc sine of 3/4

OpenStudy (anonymous):

where did you get 3/4 from ?

OpenStudy (anonymous):

first off that is wrong, so nvm secondly, unless i am miss my guess the picture is wrong too!

OpenStudy (anonymous):

how ? the 20ft is the ladder and the 15ft is the wall .

OpenStudy (anonymous):

oh maybe not fifteen feet from the floor might mean the height of the ladder

OpenStudy (anonymous):

ok fine lets say that is right then \[\sin(a)=\frac{15}{20}=\frac{3}{4}\] so \[a=\arcsin (\frac{3}{4})\]

OpenStudy (anonymous):

wow okay that makes sense . if the picture was wrong what would you do ?

OpenStudy (anonymous):

now i think it is right but if it is wrong and the 15 is along the ground, take the arc cosine

OpenStudy (anonymous):

we say that: \[\sin a=15/20 \\ \sin a =3/4 \\a=\sin^{-1}(3/4) \\ a\approx 48.6 degrees\]

OpenStudy (anonymous):

\[x=20 \cos \left(\sin ^{-1}\left(\frac{15}{20}\right)\right)=5 \sqrt{7}=13.2288 \]

OpenStudy (anonymous):

which is right ?? you all have different answers ...

OpenStudy (anonymous):

The length of the unknown triangle side is:\[\sqrt{20^2-15^2}=5 \sqrt{7}=13.2288 \]The angle is:\[\frac{180 \sin ^{-1}\left(\frac{15}{20}\right)}{\pi }=48.5904 \text{ degrees} \]

OpenStudy (anonymous):

what is the length of the third side ?

OpenStudy (anonymous):

i doubt its a decimal but i could be wrong

OpenStudy (anonymous):

never mind its right